### v1

Abstract
This paper explores the estimation and inference of the minimum spanning set (MSS), the smallest subset of risky assets that spans the mean-variance efficient frontier of the full asset set. We establish identification conditions for the MSS and develop a novel procedure for its estimation and inference. Our theoretical analysis shows that the proposed MSS estimator covers the true MSS with probability approaching 1 and converges asymptotically to the true MSS at any desired confidence level, such as 0.95 or 0.99. Monte Carlo simulations confirm the strong finite-sample performance of the MSS estimator. We apply our method to evaluate the relative importance of individual stock momentum and factor momentum strategies, along with a set of well-established stock return factors. The empirical results highlight factor momentum, along with several stock momentum and return factors, as key drivers of mean-variance efficiency. Furthermore, our analysis uncovers the sources of contribution from these factors and provides a ranking of their relative importance, offering new insights into their roles in mean-variance analysis.
Keywords: Mean-variance efficiency, Moving block bootstrap, Set inference, Spanning test.
JEL Codes: C14, C22.
1 Introduction
Conventional spanning tests, which assess whether one set of risky assets can span another, have been proposed and widely utilized in asset management and empirical research (e.g., Huberman and Kandel (1987), Ferson, Foerster, and Keim (1993), De Roon, Nijman, and Werker (2001), Amengual and Sentana (2010), Kan and Zhou (2012), Peñaranda and Sentana (2012), among others). In practice, these tests are often used to determine whether an additional set of risky assets can further extend the mean-variance efficient frontier of a given set of benchmark assets. Despite their extensive use, a notable limitation remains: to the best of our knowledge, no existing methods can estimate the smallest subset of assets that preserves the efficient frontier of the full set, including both the benchmark and the additional assets. This gap is significant given the growing demand among practitioners to identify the most relevant assets.
To address this gap, we propose an estimation procedure for identifying the minimum spanning set (MSS) within a given collection of risky assets. Formally, consider a set of $d$ assets ($d\geq 2$) represented by their returns, $R=(R_{i})_{i\leq d}$. Our objective is to assess whether the size of $R$ can be reduced without compromising its mean-variance efficiency and to identify the smallest subset of assets that reproduces the efficient frontier of the full set $R$. This subset, referred to as the MSS, is the focus of this paper.
Our research question is related to, but distinct from, those addressed by traditional spanning tests. Conventional tests evaluate whether an additional set of assets, taken as a whole, is redundant, i.e., whether adding these assets to a benchmark set extends its mean-variance efficient frontier. However, they provide no insights into the relative importance of individual assets within either the additional or benchmark set, nor do they address whether any subsets within these groups are redundant and can be excluded without compromising mean-variance efficiency.
In contrast, our method directly estimates the MSS and offers statistical insights into the relative importance of assets within the entire set. Additionally, when new assets beyond $R$ become available, our method evaluates their relevance and determines whether their inclusion renders any existing assets in $R$ redundant. This approach provides valid statistical inference on asset relevance, and is valuable for investors who aim to minimize asset management costs by identifying and investing in the smallest subset of assets capable of maintaining mean-variance efficiency.
To ensure that the estimation and inference of the MSS is well-defined, we begin by establishing its existence and uniqueness under the mild assumption that the variance-covariance matrix of $R$ is non-singular. Next, we derive the identification conditions for the MSS based on a set of restrictions on the regression coefficients. These coefficients depend exclusively on the first two moments of $R$, ensuring they are consistently estimable. Consequently, the restrictions embedded in the identification conditions are empirically testable.
We construct a statistic $M_{i,T}$, where $T$ denotes the sample size, to evaluate the identification restrictions for each asset $R_{i}$ in $R$. This statistic converges in distribution to a maximum normal distribution if $R_{i}$ is redundant and diverges to infinity if $R_{i}\in\mathrm{MSS}$. Thus, $M_{i,T}$ can be employed for a pointwise statistical inference on whether $R_{i}$ belongs to the MSS. However, since our objective is on estimation and inference of the $\mathrm{MSS}$, which is a set potentially containing multiple assets, a uniformly inference procedure based on $M_{i,T}$ over $i=1,\ldots,d$ is required.
Two technical challenges arise in conducting uniform inference. First, the (asymptotic) joint distribution of $M_{i,T}$ for $i=1,\ldots,d$ depends on unknown parameters, making it non-pivotal. To address this, we propose a resampling method based on the moving blocks bootstrap (MBB) (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) to approximate the finite-sample “null” joint distribution of $M_{i,T}$ for $i=1,\ldots,d$. The MBB also accounts for potential serial correlation in financial returns. Second, $M_{i,T}$ diverges to infinity with $T$ if and only if $i\in\mathrm{MSS}$. To ensure the inference procedure is not conservative and maintains exact control of size (Type-I error), the desired “null” joint distribution of $M_{i,T}$ should be concentrated on $i\notin\mathrm{MSS}$. However, since the $\mathrm{MSS}$ is unknown, this desired “null” joint distribution remains infeasible even with the MBB. We address this issue through a two-step approach. In the first-step, we compute a critical value based on a known upper bound of the $\mathrm{MSS}$. While conservative, this critical value ensures consistent estimation of the $\mathrm{MSS}$. In the second-step, the consistent $\mathrm{MSS}$ estimator obtained in the first-step is used to refine the critical value, enabling non-conservative and more powerful inference on the $\mathrm{MSS}$.
Our estimator of the $\mathrm{MSS}$ is formally defined as the subset of assets whose $M_{i,T}$ exceeds the refined critical value obtained through the two-step approach. Additionally, the magnitude of $M_{i,T}$ serves as a metric for evaluating the relative importance of the assets and ranking them in $R$. We theoretically demonstrate that this $\mathrm{MSS}$ estimator covers the true $\mathrm{MSS}$ with probability approaching 1 (wpa1), and converges to the exact $\mathrm{MSS}$ with probability reaching any pre-specified level, such as 0.95 or 0.99. This estimator can be made consistent by lettng the pre-specified level approach 1 with increasing sample size.
As a by-product, our $\mathrm{MSS}$ estimation procedure can also be applied to the conventional spanning test problem, offering more insights than traditional spanning tests. Given a pre-specified benchmark asset set and an additional asset set, our method can identify and estimate the $\mathrm{MSS}$ of all assets under consideration. To determine whether the benchmark set spans the additional set, we simply check whether the estimated $\mathrm{MSS}$ is a subset of the benchmark set. More importantly, by analyzing the intersections of the estimated $\mathrm{MSS}$ with the benchmark set and the additional set, we can identify which assets in the benchmark set become redundant upon including the additional set, and which assets in the additional set are truly valuable. This approach offers a more nuanced and refined assessment of asset relevance, surpassing the binary conclusions of conventional spanning tests.
The finite-sample performance of our proposed $\mathrm{MSS}$ estimation procedure is assessed through extensive Monte Carlo simulations. We simulate the data using a model with an autoregressive (AR) conditional mean and a generalized autoregressive conditional heteroskedasticity (GARCH) conditional variance, which effectively captures key stylized features of financial returns, including serial and cross-sectional correlation as well as volatility clustering. The simulation results demonstrate that the empirical probability of the estimated $\mathrm{MSS}$ containing the true $\mathrm{MSS}$ approaches one as the sample size increases. Furthermore, the empirical probability of the estimated $\mathrm{MSS}$ being identical to the true $\mathrm{MSS}$ aligns closely with the nominal significance level for sufficiently large sample sizes. These findings are consistent with the asymptotic theory estalished for our method, demonstrating its robust performance in finite samples.
We apply the proposed method to study the relative importance of stock momentum factors and factor momentum strategies, along with a set of well-established stock return factors. [^1] We thank Ehsani and Linnainmaa (2022) for kindly making their data available. The main findings from our empirical analysis are as follows. First, when either individual stock momentum factor or factor momentum is combined with the return factors, they are consistently included in the estimated $\mathrm{MSS}$, highlighting the significance of return momentum in mean-variance analysis. Second, when factor momentum coexists with all individual stock momentum factors, it is consistently selected in the $\mathrm{MSS}$. At the same time, individual stock momentum factors—such as the standard momentum and the industry-adjusted momentum—also contribute to enhancing mean-variance efficiency. Third, our empirical analysis reveals differing relative importance between the two factor momentum strategies. When both factor momentum strategies are included with other individual stock momentum factors and return factors, only the momentum in the first ten principal component factors is selected in the estimated $\mathrm{MSS}$. This result aligns with Ehsani and Linnainmaa (2022), which suggests that factor momentum effectively prices individual stock momentum and is generally concentrated in high-eigenvalue principal components. Additionally, our method underscores the importance of several individual stock momentum strategies, such as standard momentum and industry-adjusted momentum, as well as other prominent stock return factors, including excess market return, size, and betting against beta.
Our study makes a direct contribution to the growing literature on conventional spanning tests. For example, Huberman and Kandel (1987) derives the key conditions under which a given set of assets spans the mean-variance frontier of a larger set when additional assets are included, and introduces a likelihood ratio test to assess the redundancy of the additional set of assets. Subsequent advancements in this field have been made by Hansen and Jagannathan (1991), Ferson, Foerster, and Keim (1993), De Santis (1993), Bekaert and Urias (1996), De Roon, Nijman, and Werker (2001), Amengual and Sentana (2010), Kan and Zhou (2012), Peñaranda and Sentana (2012), among others. However, as emphasized earlier, our study is the first to focus on the identification and estimation of the $\mathrm{MSS}$, marking a significant departure from the existing literature on spanning tests. Empirically, our work adds to the ongoing discussions on the interplay between factor momentum and momentum factors, as detailed by Gupta and Kelly (2018), Ehsani and Linnainmaa (2022), Yan and Yu (2023), and Arnott, Kalesnik, and Linnainmaa (2023). By characterizing the $\mathrm{MSS}$ for various momentum strategies, evaluating their relative importance, and ranking them within a large set of assets, our approach provides novel insights into the interactions between these factors and their role in mean-variance analysis.
The remainder of this paper is organized as follows. Section 2 introduces the identification conditions for the MSS and details the implementation of the proposed estimation and inference method. This section also establishes the asymptotic properties of the method. Section 3 presents simulation studies to assess the finite-sample performance of the method, while Section 4 offers an empirical application. Finally, Section 5 concludes the paper. The Appendix includes proofs of the main theoretical results, auxiliary lemmas used in these proofs, and additional simulation results. The Supplemental Appendix contains detailed proofs of the auxiliary lemmas.
Notation. We use $a\equiv b$ to indicate that $a$ is defined as $b$. For any positive integer $m$, let $I_{m}$ denote the $m\times m$ identity matrix. For any positive integers $m_{1}$ and $m_{2}$, $\mathbf{1}_{m_{1}\times m_{2}}$ and $\mathbf{0}_{m_{1}\times m_{2}}$ denote the $m_{1}\times m_{2}$ matrices of ones and zeros, respectively. For real numbers $a_{1},\ldots,a_{m}$, let $(a_{i})_{i\leq m}\equiv(a_{1},\ldots,a_{m})^{\top}$, and let $a_{-i}$ denote the subvector of $(a_{i})_{i\leq m}$ with $a_{i}$ excluded. Define the support of $(a_{i})_{i\leq m}$ as $\mathrm{Supp}_{(a_{i})_{i\leq m}}\equiv\{i=1,\ldots,m:a_{i}\neq 0\}$. For any matrices $A$ and $B$, $\mathrm{diag}(A,B)$ represents a block diagonal matrix with $A$ and $B$ as its diagonal blocks, and $A\otimes B$ denotes the Kronecker product of $A$ and $B$. Additionally, $A_{j,.}$ represents the $j$th row of the matrix $A$. For any positive integer $d$, let $\mathcal{M}\equiv\{1,\ldots,d\}$, and for any positive integer $i\leq d$, let $\ell_{d,i}$ denote the $d\times 1$ vector whose $i$-th entry is $1$, with all other entries equal to 00. For two sequences of positive numbers $a_{n}$ and $b_{n}$, we write $a_{n}\succ b_{n}$ if $a_{n}\geq c_{n}b_{n}$ for some strictly positive sequence $c_{n}\rightarrow\infty$.
2 Main Theory
This section presents the main theoretical results of the paper. In subsection 2.1, we examine the existence and uniqueness of the $\mathrm{MSS}$ and establish its identification condition. The identification condition is constructive, as it is employed in subsection 2.2 to develop valid estimation and inference methods for the $\mathrm{MSS}$.
2.1 Identification Condition
For a given set of assets, represented by their returns $R\equiv(R_{i})_{i\leq d}$, our goal is to identify the smallest subset, referred to as the $\mathrm{MSS}$, such that the assets in this subset span the mean-variance frontier of the original set. The $\mathrm{MSS}$ must satisfy two key conditions: first, it must span the mean-variance frontier of the original set; second, it cannot be further reduced, meaning that any proper subset of the $\mathrm{MSS}$ cannot span the mean-variance frontier of the original set. Based on these properties, we provide the formal definition of the $\mathrm{MSS}$ below.
Definition 1.
A subvector of $R$ is called a minimum spanning set ($\mathrm{MSS}$), if it is the smallest subvector of $R$ which spans the mean-variance frontier of $R$.
Since the mean-variance frontier of $R$ depends only on its mean $\mu$ and variance $\Sigma$, which are consistently estimable, we assume in this subsection that both $\mu$ and $\Sigma$ are known for the purpose of investigating the identification condition of the $\mathrm{MSS}$. Given that the set of assets $R$ spans its own mean-variance frontier, the $\mathrm{MSS}$ is guaranteed to exist. Specifically, since there are $22^{d}-$ 2 nonempty and proper subvectors of $R$, we can examine each of these subvectors and identify those that span the mean-variance frontier of $R$. The $\mathrm{MSS}$ will be the subvector(s) in this collection with the smallest dimension. If no proper subvector spans the mean-variance frontier of $R$, then the $\mathrm{MSS}$ is $R$ itself.
A question of uniqueness naturally arises from the above discussion on the existence of the $\mathrm{MSS}$: are there two or more distinct subvectors, say $R_{K,1}$ and $R_{K,2}$ of $R$ with the same dimension, that span the mean-variance frontier of $R$? The answer is no, as demonstrated in the lemma below. This lemma also provides a constructive approach for identifying the $\mathrm{MSS}$, which serves as the basis of our proposed estimation and inference procedure.
Lemma 1.
Suppose that $\Sigma$ is finite and non-singular. Then the $\mathrm{MSS}$ exists and is unique. Moreover, for any asset $R_{i}$ in $R$, consider the least squares (LS) regression of $R_{i}$ on the remaining assets in $R$, denoted as $R_{-i}$:
$R_{i}=\alpha_{i}+\beta_{i}^{\top}R_{-i}+\varepsilon_{i}.$ (2.1)
Then the $\mathrm{MSS}$ satisfies:
$\left\{\begin{array}[]{cc}\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}\neq 0,&\text{for any }i\in\mathrm{MSS}\\ \alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}=0,&\text{for any }i\notin\mathrm{MSS}\end{array}\right..$ (2.2)
Lemma 1 has three important implications. First, each asset $R_{i}$ in $R$ can be characterized by a pair of values $\theta_{i}\equiv\left(\alpha_{i},1_{d-1}^{\top}\beta_{i}\right)^{\top}$, which are uniquely determined by $\mu$ and $\Sigma$. This pair, along with the condition in (2.2), enables us to obtain the $\mathrm{MSS}$ as follows:
$\mathrm{MSS}=\left\{i=1,\ldots,d\text{: }\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}>0\right\}.$ (2.3)
Second, the identification condition presented in Lemma 1 is constructive and does not require any prior knowledge of the $\mathrm{MSS}$. In the population where $\mu$ and $\Sigma$ of $R$ are known, we only need to compute $\theta_{i}$ for $i=1,\ldots,d$, and use (2.3) to determine the $\mathrm{MSS}$. Third, Lemma 1 strengthens a key result from Huberman and Kandel (1987), which is widely used in practice to assess the redundancy of additional assets relative to a benchmark asset set. Specifically, Proposition 3 of Huberman and Kandel (1987) implies that $R_{-i}$ spans the mean-variance frontier of $R$ (and hence $R_{i}$ is redundant relative to $R_{-i}$) if and only if:
$\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}=0.$ (2.4)
While this result identifies the redundancy of an individual asset relative to its complement in the full set, it does not provide a method for determining the $\mathrm{MSS}$. Lemma 1 extends this finding by showing that removing all assets satisfying (2.4) from $R$ yields the $\mathrm{MSS}$.
In practice, the values of $\theta_{i}$ for each asset $R_{i}$ are unknown but can be estimated through LS regression of $R_{i}$ on $R_{-i}$. Combined with the identification conditions established in Lemma 1, this enables the estimation and inference of the $\mathrm{MSS}$ in finite samples. In the next subsection, we analyze the asymptotic properties of the LS estimators of $\theta_{i}$ uniformly over $i=1,\ldots,d$. These properties facilitate the construction of $\mathrm{MSS}$ estimators that cover the true $\mathrm{MSS}$ wpa1, and asymptotically identify it with any desired level of confidence.
2.2 Estimation and Inference of the Minimum Spanning Set
We first introduce some notations to simplify the definition of the estimator for $\theta_{i}$, as well as the estimation procedure for the $\mathrm{MSS}$. The key value $\theta_{i}$ can be expressed as a linear transformation of $\alpha_{i}$ and $\beta_{i}$: $\theta_{i}=A\cdot(\alpha_{i},\beta_{i}^{\top})^{\top}$ where $A\equiv\mathrm{diag}(1,1_{d-1}^{\top})$. For any observation $R_{t}\equiv(R_{i,t})_{i\leq d}$ where $t=1,\ldots T$, we let $\tilde{R}_{-i,t}\equiv[1,R_{-i,t}^{\top}]^{\top}$ denote the regressors in the linear regression specified in (2.1), $\hat{Q}_{-i}\equiv T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}$ and $Q_{-i}\equiv\mathrm{E}[\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}]$ for any $i=1,\ldots,d$.
Next, we describe the estimation procedures for the $\mathrm{MSS}$ and their intuition. For each asset $R_{i}$, the LS estimators of $\alpha_{i}$ and $\beta_{i}$ are defined as:
$(\hat{\alpha}_{i},\hat{\beta}_{i}^{\top})^{\top}\equiv\hat{Q}_{-i}T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}R_{i,t}.$ (2.5)
Given the definition of $\theta_{i}$, we can estimate it with $\hat{\theta}_{i}\equiv A\cdot(\hat{\alpha}_{i},\hat{\beta}_{i}^{\top})^{\top}$. Using the expression for $R_{i}$ in (2.1), the definition of $\hat{\theta}_{i}$ and the LS estimators  in (2.5), we obtain an expression for the estimation error in $\hat{\theta}_{i}$:
$T^{1/2}(\hat{\theta}_{i}-\theta_{i})=A\hat{Q}_{-i}^{-1}\left(T^{-1/2}\sum_{t\leq T}\tilde{R}_{-i,t}\varepsilon_{i,t}\right).$ (2.6)
Let $\Omega_{d,T,i}\equiv\mathrm{Var}(T^{-1/2}\sum_{t\leq T}\tilde{R}_{-i,t}\varepsilon_{i,t})$. For any fixed $i$, we can use this expression, and apply the law of large numbers (LLN) and the central limit theorem (CLT) to show that $T^{1/2}(\hat{\theta}_{i}-\theta_{i})$ is approximately distributed as normal with mean zero and variance $AQ_{-i}^{-1}\Omega_{d,T,i}Q_{-i}^{-1}A^{\top}$, denoted as
[…]
This pointwise result can be used to test whether a given asset $R_{i}$ belongs to the $\mathrm{MSS}$ or not. However, our goal is to estimate and conduct statistical inference on the $\mathrm{MSS}$, which may include multiple assets. Therefore, to ensure accurate estimation of the $\mathrm{MSS}$ and proper control of statistical inference errors, we need to conduct a joint statistical inference on condition (2.2) for $i=1,\ldots,d$, which requires approximating the finite-sample distribution of $\hat{\theta}_{i}$ uniformly over $i=1,\ldots,d$.
For the purpose of joint inference, we stack the expression in (2.6) for different $i$ to obtain a joint representation of the estimation errors for $\hat{\theta}$:
$T^{1/2}(\hat{\theta}-\theta)=\left(A\hat{Q}_{-i}^{-1}T^{-1/2}\sum_{t\leq T}\tilde{R}_{-i,t}\varepsilon_{i,t}\right)_{i\leq d},$ (2.7)
where $\hat{\theta}\equiv(\hat{\theta}_{i})_{i\leq d}$ and $\theta\equiv(\theta_{i})_{i\leq d}$. By the (uniform) consistency of $\hat{Q}_{-i}$, the term on the right hand side of (2.7) can be approximated by $T^{-1/2}\sum_{t\leq T}e_{t}$, where $e_{t}\equiv(e_{i,t})_{i\leq d}$ and $e_{i,t}\equiv AQ_{-i}^{-1}\tilde{R}_{-i,t}\varepsilon_{i,t}$. Intuitively, by a CLT-type of argument, the finite sample distribution of $T^{-1/2}\sum_{t\leq T}e_{t}$ can be approximated by $\Omega_{d,T}^{1/2}\mathcal{N}_{d}$, where $\Omega_{d,T}\equiv\mathrm{Var}(T^{-1/2}\sum_{t\leq T}e_{t})$ and $\mathcal{N}_{d}$ denotes a standard normal random vector. Therefore, the finite-sample distribution of $T^{1/2}(\hat{\theta}-\theta)$ can be approximated by the distribution of $\Omega_{d,T}^{1/2}\mathcal{N}_{d}$, denoted as
[…]
This intuition is employed to obtain critical values in our procedure for estimating the $\mathrm{MSS}$.
We are now ready to introduce the test statistic and critical value used in our estimation procedure. Let $A_{j,\cdot}$ denote the $j$th row of $A$ for $j=1,2$. For each asset $R_{i}$, we construct
$M_{i,T}\equiv\max\left\{\frac{T^{1/2}\left|\hat{\alpha}_{i}\right|}{\hat{s}_{i,1}},\text{ \ }\frac{T^{1/2}\left|1_{d-1}^{\top}\hat{\beta}_{i}-1\right|}{\hat{s}_{i,2}}\right\},$ (2.9)
where $\hat{s}_{i,j}^{2}\equiv\hat{\sigma}_{\varepsilon_{i}}^{2}A_{j,\cdot}\hat{Q}_{-i}^{-1}A_{j,\cdot}^{\top}$ and $\hat{\sigma}_{\varepsilon_{i}}^{2}\equiv T^{-1}\sum_{t\leq T}(R_{i,t}-\hat{\theta}_{i}^{\top}\tilde{R}_{-i,t})^{2}$ for $j=1,2$. Clearly, $M_{i,T}$ is the maximum of the t-ratios for testing $\alpha_{i}=0$ and $11_{d-1}^{\top}\beta_{i}=$ 1, respectively. Given the identification condition in (2.2) and the approximation result in (2.8), it follows that for any $i\in\mathrm{MSS}$, $M_{i,T}$diverges as the sample size $T$ increases, while for any $i\notin\mathrm{MSS}$$M_{i,T}$ can be approximated in distribution by
$\tilde{M}_{i,T}\equiv\left\|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d,T}^{1/2}\mathcal{N}_{d}\right\|_{\infty},$ (2.10)
where $s_{i,j}^{2}\equiv\sigma_{\varepsilon_{i}}^{2}A_{j,\cdot}Q_{-i}^{-1}A_{j,\cdot}^{\top}$ and $\sigma_{\varepsilon_{i}}^{2}\equiv\mathrm{E}[\varepsilon_{i,t}^{2}]$. This implies that the assets in the $\mathrm{MSS}$ tend to have larger $M_{i,T}$ values than those that are not in the $\mathrm{MSS}$. A formal statistical inference procedure should provide a critical value under a pre-specified significance level $p\in(0,1)$ to determine when an asset with a large $M_{i,T}$ value can be included in the $\mathrm{MSS}$.
Ideally, for any small $p\in(0,1)$, the critical value should depend only on the assets that are not in the $\mathrm{MSS}$. This is because, in view of Lemma 1 the “null hypothesis”:
$\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}=0$
holds only for $i\notin\mathrm{MSS}$. Therefore, if the $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, we would use the $(1-p)$-quantile of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, denoted as $cv_{1-p}^{u}$, as the critical value. On the other hand, if $\mathrm{MSS}=\mathcal{M}$, we shall set $cv_{1-p}^{u}$ to any fixed constant. From the discussion in the previous paragraph, as $T$approaches infinity, we have
$\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p}^{u}\right)\rightarrow 1$ (2.11)
regardless of whether the $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, and
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}>cv_{1-p}^{u}\right)\rightarrow p,$ (2.12)
if $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$. This implies that the estimator of $\mathrm{MSS}$ defined as:
$\widehat{\mathrm{MSS}}_{p}^{u}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>cv_{1-p}^{u}\right\}$ (2.13)
has the desirable properties that it covers the true $\mathrm{MSS}$ wpa1, and overestimates it with a small probability $p$.
However, the critical value $cv_{1-p}^{u}$ is not practical to use for two main reasons: (i) the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ is unknown due to the presence of nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$; and (ii) there is no prior knowledge about the $\mathrm{MSS}$. The first challenge can be addressed by estimating the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ through either plugging in consistent estimators of the nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$ directly, or by using resampling methods such as the bootstrap. In this paper, we employ the moving blocks (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) bootstrap (MBB) to approximate the distribution of $(\tilde{M}_{i,T})_{i\leq d}$.
The second challenge is more delicate. One solution is to use a known upper bound of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, e.g., $\max_{i\leq d}\tilde{M}_{i,T}$, to obtain a critical value which is larger than $cv_{1-p}^{u}$. While straightforward to implement, this method is conservative, potentially leading to an underestimation of the $\mathrm{MSS}$ in finite samples. An alternative approach involves replacing the $\mathrm{MSS}$ in $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$ with a preliminary consistent estimator, which can be obtained using the first method by letting $p$ approach zero. In this paper, we study both methods and refer to their corresponding $\mathrm{MSS}$ estimators as the one-step estimator and the two-step estimator, respectively.
Let $\hat{\alpha}_{i}^{b}$, $\hat{\beta}_{i}^{b}$ and $\hat{s}_{i,j}^{b}$ denote the bootstrap counterparts of $\hat{\alpha}_{i}$, $\hat{\beta}_{i}$ and $\hat{s}_{i,j}$, respectively. [^2] The details of the MBB bootstrap are provided in the algorithm below. For $i\leq d$, define:
$M_{i,T}^{b}\equiv\max\left\{\frac{T^{1/2}\left|\hat{\alpha}_{i}^{b}-\hat{\alpha}_{i}\right|}{\hat{s}_{i,1}^{b}},\text{\ }\frac{T^{1/2}\left|1_{d-1}^{\top}(\hat{\beta}_{i}^{b}-\hat{\beta}_{i})\right|}{\hat{s}_{i,2}^{b}}\right\},$ (2.14)
which represents the bootstrap version of $M_{i,T}$. We use the $(1-p)$-quantile of $\max_{i\leq d}M_{i,T}^{b}$, denoted as $\overline{cv}_{1-p,T}^{b}$, as the critical value to construct the one-step estimator of the $\mathrm{MSS}$:
$\overline{\mathrm{MSS}}_{p}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>\overline{cv}_{1-p,T}^{b}\right\}.$ (2.15)
Since the $\mathrm{MSS}$ is a subset of $\mathcal{M}$, the critical value $\overline{cv}_{1-p,T}^{b}$ provides an upper bound for the infeasible critical value $cv_{1-p,T}^{u}$. This ensures that the probability of $\overline{\mathrm{MSS}}_{p}$ overestimating the $\mathrm{MSS}$ is less than $p$. The properties of the one-step estimator $\overline{\mathrm{MSS}}_{p}$ are formally presented in the theorem below.
Theorem 1.
Suppose that Assumptions  1 and 2 in the Appendix holds, and moreover
$\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)\succ T^{-1}.$ (2.16)
Then we have $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}\subset\overline{\mathrm{MSS}}_{p})=1$ and $\lim\inf_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\overline{\mathrm{MSS}}_{p})\geq 1-p$.
Theorem 1 shows that the one-step estimator $\overline{\mathrm{MSS}}_{p}$ will not underestimate the $\mathrm{MSS}$, meaning that all assets in the $\mathrm{MSS}$ are included in $\overline{\mathrm{MSS}}_{p}$ wpa1. Additionally, as the sample size increases, $\overline{\mathrm{MSS}}_{p}$ identifies the $\mathrm{MSS}$ with a probability greater than $1-p$. Therefore, the probability that $\overline{\mathrm{MSS}}_{p}$ overestimates the $\mathrm{MSS}$ is asymptotically less than $p$.
In practice, one can choose a small value for $p$, such as 0.05 or 0.01, to ensure that $\overline{\mathrm{MSS}}_{p}$ asymptotically identifies the $\mathrm{MSS}$ with a high probability, such as 0.95 or 0.99. Alternatively, one may set $p=\gamma_{T}$, where $\gamma_{T}$ approaches zero as the sample size increases, ensuring that $\overline{\mathrm{MSS}}_{p}$ identifies the $\mathrm{MSS}$ consistently. Indeed, as we will show below, this is the approach we adopt in the first step of our two-step estimation procedure of the $\mathrm{MSS}$.
The condition in (2.16) specifies the least favorable scenario under which the one-step estimator $\overline{\mathrm{MSS}}_{p}$ will still cover the $\mathrm{MSS}$ wpa1 and overestimates it with probability asymptotically less than $p$. This condition is derived from the identification conditions in (2.2), and is similarly adjusted to resemble the local power analysis of statistical hypothesis test. If there exists a constant $K>0$ such that (2.2) holds for any asset $R_{i}$ in $\mathrm{MSS}$ with
$\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}>K,$
then (2.16) is trivially satisfied. More importantly, (2.16) allows for the identification conditions (2.2) to nearly hold for some assets, in the sense that their values of $\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}$ are close to zero.
Compared to the desirable estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, which is based on the infeasible critical value $cv_{1-p}^{u}$, it is evident that $\overline{\mathrm{MSS}}_{p}\subset\widehat{\mathrm{MSS}}_{p}^{u}$ with high probability because $cv_{1-p}^{u}\leq\overline{cv}_{1-p,T}^{b}$ wpa1. This implies that $\overline{\mathrm{MSS}}_{p}$ has a higher probability of missing assets in the $\mathrm{MSS}$ compared to $\widehat{\mathrm{MSS}}_{p}^{u}$ in finite samples, although asymptotically, both should include the $\mathrm{MSS}$ wpa1. The reason for this unfavorable property of the one-step estimator $\overline{\mathrm{MSS}}_{p}$ is that its critical value $\overline{cv}_{1-p,T}^{b}$ is asymptotically larger than the value $cv_{1-p}^{u}$ required for exact control of the type-I error (i.e., the error of overestimating the $\mathrm{MSS}$) in statistical inference. This leads to potential power loss in detecting assets in the $\mathrm{MSS}$, especially when the size of the $\mathrm{MSS}$ is relatively large.
To mitigate the power loss of the one-step estimator $\overline{\mathrm{MSS}}_{p}$, we now introduce the two-step procedure for estimating the $\mathrm{MSS}$. In the first step, we obtain a preliminary estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ for some positive sequence $\gamma_{T}=o(1)$. [^3] We recommend $\gamma_{T}=\min\{p/2,$ $0.05/\log(\log(T))\}$, which is used in the simulation studies and empirical application of this paper. In view of Theorem 1 with $p=\gamma_{T}$, the preliminary estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ should be identical to the $\mathrm{MSS}$ wpa1. This motivates us to estimate the infeasible critical value $cv_{1-p}^{u}$, the $(1-p)$-quantile of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, by the $(1-p)$-quantile of $\max_{i\notin\overline{\mathrm{MSS}}_{\gamma_{T}}}M_{i,T}^{b}$. The latter is denoted as $cv_{1-p,T}^{b}$. When $\overline{\mathrm{MSS}}_{\gamma_{T}}=\mathcal{M}$, we set $cv_{1-p,T}^{b}=0$. Subsequently, our estimate of the $\mathrm{MSS}$ is:
$\widehat{\mathrm{MSS}}_{p}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>cv_{1-p,T}^{b}\right\}.$ (2.17)
The implementation details of the two-step procedure are provided in the algorithm below.
From the definitions of $M_{i,T}$ and $\widehat{\mathrm{MSS}}_{p}$, we have:
$\widehat{\mathrm{MSS}}_{p}=\left\{i\leq d\text{: \ }\left|\hat{\alpha}_{i}\right|>T^{-1/2}\hat{s}_{i,1}cv_{1-p,T}^{b}\right\}\bigcup\left\{i\leq d\text{: }|1_{d-1}^{\top}\hat{\beta}_{i}-1|>T^{-1/2}\hat{s}_{i,2}cv_{1-p,T}^{b}\right\}.$ (2.18)
Here, $(T^{-1/2}\hat{s}_{i,1}cv_{1-p,T}^{b})_{i\leq d}$ and $(T^{-1/2}\hat{s}_{i,2}cv_{1-p,T}^{b})_{i\leq d}$ can be interpreted as uniform confidence bands for $\left|\alpha_{i}\right|$ and $|1_{d-1}^{\top}\beta_{i}-1|$, respectively. From the expression in (2.18), it follows that an asset $i$ is selected into $\widehat{\mathrm{MSS}}_{p}$, if either $\left|\hat{\alpha}_{i}\right|$, $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both, exceed their respective uniform confidence bands. [^4] As noted in Kan and Zhou (2012), these two cases, i.e., $|\alpha_{i}|>0$ and $|1_{d-1}^{\top}\beta_{i}-1|$, are subject to explicit economic interpretation. Specifically, $|\alpha_{i}|>0$ implies asset $i$ contributes to the tangency portfolio, while $|1_{d-1}^{\top}\beta_{i}-1|>0$ corresponds to the scenario where asset $i$ contributes to the global minimum variance portfolio. For more details on the derivations, see Section 2.1 in Kan and Zhou (2012). Therefore, our method provides detailed insights into the sources of contribution for the selected assets. Additionally, by ranking all assets in the full set according to their $M_{i,T}$ values, we can directly measure their relative importance. [^5] For an empirical illustration of these ideas, refer to Figures 2 and 3, , as well as the related discussion in Section 4.
Algorithm: A Bootstrap $\mathbf{MSS}$ Estimation Procedure
Step 1. For each asset $R_{i}$, run the linear regression specified in (2.1) to obtain the estimators of $\hat{\alpha}_{i}$ and $1_{d-1}^{\top}\hat{\beta}_{i}^{\top}$, and calculate $M_{i,T}$ specified in (2.9).
Step 2. Given a bandwidth $\ell$, define moving blocks $B_{j}=\{R_{j},\ldots,R_{j+\ell-1}\}$ for $j=1,\ldots,q$ where $q\equiv T-\ell+1$. This will give $q$ blocks $\{B_{j}\}_{j\leq q}$.
Step 3. Let $m\equiv[T/\ell]$ and $T_{B}\equiv m\ell$. Resample $m$ blocks $Z_{j}^{b}\equiv\{R_{\ell(j-1)+1}^{b},\ldots,R_{\ell(j-1)+\ell}^{b}\}$ independently from $\{B_{j}\}_{j\leq q}$ to obtain a bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}\equiv\{Z_{j}^{b}\}_{j\leq m}$, and then calculate $M_{i,T}^{b}$ specified in (2.14).
Step 4. Repeat Step 3 $B$ times, and set $\overline{cv}_{1-\gamma_{T},T}^{b}$ to the $(1-\gamma_{T})$-sample quantile of $\{\max_{i\leq d}M_{i,T}^{b}\}_{b\leq B}$.
Step 5. Find the preliminary $\mathrm{MSS}$ estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ using $\overline{cv}_{1-\gamma_{T},T}^{b}$ and the formula in (2.15).
Step 6. If $\overline{\mathrm{MSS}}_{\gamma_{T}}=\mathcal{M}$, set $cv_{1-p,T}^{b}=0$. Otherwise set $cv_{1-p,T}^{b}$ to the $(1-p)$-sample quantile of $\{\max_{i\notin\overline{\mathrm{MSS}}_{\gamma_{T}}}M_{i,T}^{b}\}_{b\leq B}$, and obtain the $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ using the formula in (2.17).
Step 7. If $\widehat{\mathrm{MSS}}_{p}$ is an empty set in Step 6, set $\widehat{\mathrm{MSS}}_{p}=\{i=1,\ldots,d:M_{i,T}\geq\max_{j\leq d}M_{j,T}\}$.
The last step in the algorithm above is necessary as a finite sample adjustment. When the size of the $\mathrm{MSS}$ is small and the identification condition in (2.2) nearly holds, the test based on the estimated critical value $cv_{1-p,T}^{b}$ may cause the estimator $\widehat{\mathrm{MSS}}_{p}$ to be an empty set with a small sample size. However, since the $\mathrm{MSS}$ cannot be empty, we estimate it as the set of assets that have the highest values of $M_{i,T}$ in this case.
The above algorithm can be simplified to obtain an one-step estimator $\overline{\mathrm{MSS}}_{p}$ that satisfies Theorem 1. Specifically, in Steps 4 and 5 of the algorithm, we replace $\gamma_{T}$ with $p$ to obtain $\overline{\mathrm{MSS}}_{p}$. If $\overline{\mathrm{MSS}}_{p}$ results in an empty set, we redefine it according to Step 7 of the algorithm.
Theorem 2.
Suppose that Assumptions  1 and 2 in the Appendix holds, and moreover
$\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)\succ\gamma_{T}^{-1}T^{-1}.$ (2.19)
Then we have $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}}_{p})=1$. Moreover, if $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, then $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1-p$.
Theorem 2 establishes that the two-step $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ possesses the same desirable properties as the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, as conjectured in (2.11) and (2.12). Specifically, $\widehat{\mathrm{MSS}}_{p}$ covers the $\mathrm{MSS}$ wpa1, and overestimates the $\mathrm{MSS}$ with probability exactly $p$ as the sample size $T$ increases. If $\mathrm{MSS}=\mathcal{M}$, then the first result of Theorem 2 implies that $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1$.
Theorem 2 also shows that the two-step $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ shares similar properties with $\widehat{\mathrm{MSS}}_{p}^{u}$ when compared to the one-step estimator $\overline{\mathrm{MSS}}_{p}$. Since $cv_{1-p,T}^{b}\leq\overline{cv}_{1-p,T}^{b}$ by construction, the one-step estimator $\overline{\mathrm{MSS}}_{p}$ is always a subset of the two-step estimator $\widehat{\mathrm{MSS}}_{p}$ in finite samples, although both are expected to cover the $\mathrm{MSS}$ wpa1. This implies that, similar to the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, the probability that $\widehat{\mathrm{MSS}}_{p}$ misses any assets in the $\mathrm{MSS}$ is smaller than that of $\overline{\mathrm{MSS}}_{p}$.
3 Monte Carlo Simulations
In this section, we examine the finite-sample performance of the proposed MSS estimators using Monte Carlo simulations. The simulation design is detailed in subsection 3.1, and the results are presented in subsection 3.2.
3.1 The Simulation Setting
We use a vector autoregressive (VAR) model to specify the conditional mean and a GARCH model to capture the conditional variance of returns. This VAR-GARCH framework effectively reproduces key stylized features of stock returns, including serial and cross-sectional correlations as well as volatility clustering.
Specifically, the returns are generated using the following equations:
$R_{K,t}=\mu\cdot\mathbf{1}_{K}+\mathbf{A}R_{K,t-1}+\eta_{K,t},$ (3.1)
and
$R_{N,t}=\mathbf{a}+\mathbf{B}R_{K,t}+\eta_{N,t}.$ (3.2)
Here, (3.1) specifies a VAR of order one for the assets $R_{K,t}$ in the benchmark MSS, where $\mu\cdot\mathbf{1}_{K}$ and $\mathbf{A}$ denote the mean vector and the autoregressive coefficient matrix, respectively. The returns of possible redundant assets $R_{N,t}$ are connected to those in the benchmark MSS via (3.2), where $\mathbf{a}$ and $\mathbf{B}$ are parameters used to define the actual MSS in accordance with Lemma 1.
To incorporate GARCH effects, we define $\eta_{t}\equiv(\eta_{K,t}^{\top},\eta_{N,t}^{\top})^{\top}$, where the $i$th component of $\eta_{t}$, denoted as $\eta_{i,t}$ for $i=1,\ldots,K+N$, satisfies:
$\eta_{i,t}$ $=$ $d_{i,t}v_{i,t},$
$d_{i,t}$ $=$ $(0.1+0.1\eta_{i,t-1}^{2}+0.8d_{i,t-1}^{2})^{1/2},$
with $(v_{i,t})_{i\leq K+N}$ being a $(K+N)\times 1$ standard normal random vector.
In the simulation, we consider $K+N=8$, where $K\in\{1,3,5,7\}$, and the sample size $T\in\{120,180,240,300\}$, corresponding to monthly observations spanning 10 to 25 years—periods commonly encountered in practice. For the data-generating process of the MSS asset returns, we set $\mu=0$ and define the autoregressive matrix $\mathbf{A}\equiv(a_{i,j})_{i,j\leq K}$ with $a_{i,j}=\rho^{|i-j|+1}$ and $\rho=0.1$. We ensure the assets whose returns are governed by (3.2) are redundant by setting $\mathbf{a}=\mathbf{0}_{N\times 1}$ and constraining the row sums of $\mathbf{B}$ to equal 1. For simplicity, we specify $\mathbf{B}=K^{-1}\mathbf{1}_{N\times K}$. The block size $\ell$ in the MBB procedure is set to $\lfloor 1.2T^{1/4}\rfloor$.
Three MSS estimators are investigated in the simulation study: the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, the one-step estimator $\overline{\mathrm{MSS}}_{p}$, and the two-step estimator $\widehat{\mathrm{MSS}}_{p}$, as defined in (2.13), (2.15) and (2.17), respectively. For these estimators, we set $p=0.05$, and their finite-sample performance is assessed using 10,000 Monte Carlo replications.
3.2 Simulation Results
Guided by theoretical insights, we first examine the empirical probabilities of the estimated MSS containing the true MSS, as reported in the second to fourth columns of Table 1. Consistent with the main theory established in Section 2, the simulation results indicate that both the one-step and two-step MSS estimators include the true MSS with high probability across all combinations of $K$ and $N$, provided the sample size is moderately large (e.g., $T\geq 180$). This probability approaches 1 for the two-step estimator when the sample size increases to 240, and for the one-step estimator when the sample size reaches approximately $300$.
Compared with the one-step estimator, the two-step MSS estimator exhibits a higher probability of containing the true MSS, especially when $N$ (the size of redundant set) is small. As discussed in Section 2, this is because the critical value employed by the two-step estimator is no larger than that of the one-step estimator, and their difference becomes more pronounced as $N$ decreases. The infeasible MSS estimator is the one which the two-step MSS estimator tries to mimic. As we can see from Table 1, their properties become identical for all combinations of $K$ and $N$ when the sample size becomes large. On the other hand, the one-step estimator is almost identical to the infeasible estimator when $K=1$, but it is more likely to miss assets in the true MSS when both $N$ and $T$ are small.
Next, we evaluate the ability of the estimated MSS to identify the true MSS, measured as the probability that the MSS estimator exactly matches the true MSS. This property is particularly important in practice, as it demonstrates the effectiveness of the MSS estimators in reducing the number of assets while maintaining the mean-variance efficient frontier. The results, presented in the fifth to seventh columns of Table 1, show that both the one-step and two-step MSS estimators accurately identify the true MSS with high probability when the sample size is sufficiently large, across all combinations of $K$ and $N$ (e.g., $T\geq 180$). When the sample size $T$ reaches $240$, the probabilities of the two-step MSS estimator coinciding with the true MSS exceed $0.9$ and converge to the nominal value of $0.95$ as $T=300$.
Table 1: Simulation Results for the MSS Estiamtors
$\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}})$ $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}})$
One-step Est. Two-step Est. Infeasible Est. One-step Est. Two-step Est. Infeasible Est.
Panel A: $(K,N)=(1,7)$
$T=120$ 0.939 0.939 0.939 0.911 0.909 0.908
$T=180$ 0.986 0.986 0.987 0.953 0.948 0.948
$T=240$ 0.998 0.998 0.998 0.959 0.954 0.954
$T=300$ 0.999 0.999 0.999 0.956 0.950 0.950
Panel B: $(K,N)=(3,5)$
$T=120$ 0.898 0.913 0.923 0.873 0.872 0.881
$T=180$ 0.992 0.994 0.996 0.964 0.951 0.952
$T=240$ 0.999 0.999 0.999 0.970 0.953 0.953
$T=300$ 1.000 1.000 1.000 0.971 0.954 0.955
Panel C: $(K,N)=(5,3)$
$T=120$ 0.706 0.790 0.826 0.695 0.757 0.789
$T=180$ 0.966 0.981 0.986 0.948 0.936 0.939
$T=240$ 0.998 0.999 0.999 0.979 0.952 0.952
$T=300$ 1.000 1.000 1.000 0.982 0.948 0.948
Panel D: $(K,N)=(7,1)$
$T=120$ 0.266 0.512 0.722 0.264 0.493 0.686
$T=180$ 0.819 0.931 0.967 0.815 0.887 0.917
$T=240$ 0.969 0.992 0.996 0.964 0.943 0.946
$T=300$ 0.997 0.999 1.000 0.992 0.945 0.946
  • Note: This table presents the simulation results for the MSS estimators described in Section 2. The second to fourth columns display the empirical probabilities of the estimated MSS containing the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}})$, for the one-step, two-step, and infeasible MSS estimators. The fifth to seventh columns report the empirical probabilities of correctly identifying the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}})$, for the one-step, two-step, and infeasible MSS estimators, respectively. The significance level $p$ is set at 0.05. All empirical probabilities are calculated based on 10,000 Monte Carlo replications.
In contrast, the one-step MSS estimator is more aggressive in reducing the size of the estimated set. Its probability of identifying the true MSS is lower than that of the two-step estimator when both $N$ and $T$ are small, but surpasses the two-step estimator’s probability as $T$ increases. This higher probability for the one-step MSS estimator at large $T$ arises due to its failure to maintain exact size control. The two-step MSS estimator can also achieve a probability of identifying the true MSS greater than 0.95 with a large sample size by specifying a smaller significance level $p$, such as $p=0.01$. However, the trade-off is that this adjustment increases the likelihood of missing assets from the true MSS when the sample size is small, resembling the properties of the one-step estimator.
Figure 1: Empirical Distribution of the Size of the Estimated MSS
Note: This figure illustrates the empirical distribution of the size of the estimated MSS for four different settings. The upper panel corresponds to the settings of $(K,N,T)=(1,7,120)$ and $(K,N,T)=(1,7,240)$, respectively, while the lower panel corresponds to the settings of $(K,N,T)=(7,1,120)$ and $(K,N,T)=(7,1,240)$, respectively. The significance level $p$ is set at 0.05. All empirical probabilities are calculated based on 10,000 Monte Carlo replications.
To further investigate the size of the estimated MSS, we present the empirical distributions of the sizes of the one-step and two-step MSS estimators in Figure 1 for $(K,N)=(1,7)$ and $(7,1)$ at $T=120$ and $240$. The results indicate that the empirical distribution of the estimated MSS size is nearly identical for both estimators when the size of the redundant set is large i.e., $N=7$. However, as shown in the lower-left panel of Figure 1, the one-step estimator may exclude assets belonging to the true MSS, leading to an underestimation of its size when both $N$ and $T$ are small. [^6] Additionally, we present the empirical distribution of the size of the intersection between the estimated MSS and the true MSS (i.e., $\cap\text{MSS}$ MSS) in Figure 4 in the Appendix. These results further demonstrate that the two-step test generally performs better when $N$ is small.
In summary, the simulation results in Table 1 and Figure 1 align with our theoretical findings: when the sample size is sufficiently large, the estimated MSS not only contains but also coincides with the true MSS with high probability. These results strongly support the practical effectiveness of the proposed MSS estimators, particularly the two-step estimator, in real-world applications.
4 An Empirical Application
Momentum, which posits that assets’ past returns can effectively predict their future returns, has been identified across various asset classes and time periods. Due to its prevalence, momentum has been incorporated into widely used asset pricing models to explain cross-sectional returns, such as the Carhart four-factor model (Carhart, 1997). Conversely, several theories have been proposed to explain this anomaly, including time-varying risk , behavioral biases, and trading frictions; see Jegadeesh and Titman (2011) for a review of the momentum literature. Interestingly, Ehsani and Linnainmaa (2022) recently demonstrated that momentum in individual stock returns can be attributed to momentum in other risk factors. This finding challenges the widely accepted view in the literature that momentum is an independent risk factor.
One of the key findings in Ehsani and Linnainmaa (2022) is that momentum in factor returns effectively explains various forms of individual stock momentum, including standard stock return momentum, industry momentum, industry-adjusted momentum, intermediate momentum, and Sharpe ratio momentum. Specifically, when regressing the monthly returns of individual stock momentum strategies on the Fama-French five factors and factor momentum, the authors demonstrate that none of the alphas in the asset pricing model are statistically significant. Conversely, individual stock factors and the Fama-French five factors fail to account for the abnormal returns of factor momentum. [^7] For more details, see Table 5 and the corresponding discussions in Ehsani and Linnainmaa (2022).
In this empirical application, we further shed light on the interaction between individual stock momentum and factor momentum by estimating the $\mathrm{MSS}$ for various combinations of momentum strategies. We address the following two research questions. First, do the individual stock momentum factor and the factor momentum independently contribute to the mean-variance efficiency when analyzed alongside other well-established factors? Second, what role does factor momentum play in constructing the mean-variance efficient frontier when it coexists with the individual stock momentum factor? In contrast to Ehsani and Linnainmaa (2022), our analysis emphasizes the relative importance of various momentum strategies in shaping the mean-variance efficient frontier within the corresponding factor portfolios. [^8] For the sake of brevity, we focus on the estimation results of the two-step MSS test conducted at a significance level of 0.05. If a momentum strategy is more likely to be included in the $\mathrm{MSS}$, it should be regarded as a critical component in mean-variance analysis.
4.1 Data
Our empirical application utilizes the same dataset of monthly factor returns in the U.S. stock market as employed in Ehsani and Linnainmaa (2022). As noted in the original study, the factor data are sourced from three primary providers: Kenneth French’s website, AQR, and Robert Stambaugh. [^9] These factor returns are available for download at http://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html, https://www.aqr.com/insights/datasets, and https://finance.wharton.upenn.edu/~stambaug/. When not directly available, factor returns are computed as the difference between the average returns of the top three deciles and the bottom three deciles. The construction of these decile portfolios strictly follows the methodology outlined in the corresponding reference.
We study the $\mathrm{MSS}$ for five individual stock momentum factors and two factor momentum strategies analyzed in Ehsani and Linnainmaa (2022). The five individual stock momentum factors include the standard individual stock momentum of Jegadeesh and Titman (1993), the industry-adjusted momentum of Cohen and Polk (1998), the industry momentum of Moskowitz and Grinblatt (1999), the intermediate momentum of Novy-Marx (2012), and the Sharpe ratio momentum of Rachev, Jašić, Stoyanov, and Fabozzi (2007). [^10] These individual stock momentum factors are constructed using the common up-minus-down method with return-sorted portfolios. Specifically, the standard individual stock momentum sorts stocks based on prior returns over the past four quarters, the industry-adjusted momentum sorts stocks by prior industry-adjusted returns, intermediate momentum sorts stocks by returns from month $t-12$ to $t-7$, the Sharpe ratio momentum sorts stocks by returns scaled by return volatility, and industry momentum sorts 20 industries based on their prior six-month returns. The two factor momentum strategies are the time-series momentum applied to 20 “off-the-shelf” individual factors and the momentum in the first 10 principal component (PC) factors extracted from the 47 factors studied in Kozak, Nagel, and Santosh (2020). For further details on the construction of these factor momentum strategies, refer to Ehsani and Linnainmaa (2022). Panel A of Table 2 presents references and summary statistics for these seven momentum-related factors.
In addition, we consider a comprehensive set of well-established factors in our analysis of the $\mathrm{MSS}$ of momentum strategies, including excess market return, size, value, profitability, investment, accruals, betting against beta, cash flow to price, earnings to price, liquidity, long-term reversals, net share issues, quality minus junk, residual variance, and short-term reversal. Panel B of Table 2 provides the corresponding references and summary statistics for these factors.
Turning to the descriptive statistics reported in Table 2, we observe significant variations in the return performance of both momentum strategies and other factors. For example, for the two factor momentum strategies (FTM and PCM), the momentum in individual factors demonstrates a substantially higher monthly return (0.33% vs. 0.19%) and volatility (1.20% vs. 0.64%) compared to the momentum in the first 10 principal component factors, whereas the average monthly returns of individual stock factors are generally higher than those of factor momentum strategies, albeit with significantly greater monthly volatility. Meanwhile, the betting-against-beta factor generates the highest monthly return of 0.88% with a monthly volatility of 3.34%.
Table 2: Details on Stock Momentums, Factor Momentums, and other Factors
Factors Acronym Reference Sample Period Mean S.D.
Panel A: Momentum Factors
Momentum in individual factors FTM Ehsani and Linnainmaa (2022) 1964/07-2019/12 0.33% 1.20%
Momentum in PC factors 1-10 PCM Ehsani and Linnainmaa (2022) 1973/07-2019/12 0.19% 0.64%
Standard stock momentum MOM Jegadeesh and Titman (1993) 1964/07-2019/12 0.64% 4.22%
Industry-adjusted momentum IAM Cohen and Polk (1998) 1964/07-2019/12 0.41% 2.64%
Industry momentum IDM Moskowitz and Grinblatt (1999) 1964/07-2019/12 0.63% 4.60%
Intermediate momentum ITM Novy-Marx (2012) 1964/07-2019/12 0.48% 3.02%
Sharpe ratio momentum SRM Rachev, Jašić, Stoyanov, and Fabozzi (2007) 1964/07-2019/12 0.55% 3.59%
Panel B: Other Factors
Excess market return MKT Sharpe (1964) 1964/07-2019/12 0.53% 4.42%
Size SMB Banz (1981) 1964/07-2019/12 0.16% 3.01%
Value HML Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.31% 2.88%
Profitability RMW Novy-Marx (2013) 1964/07-2019/12 0.27% 2.21%
Investment CMA Titman, Wei, and Xie (2004) 1964/07-2019/12 0.32% 2.00%
Accruals ACC Sloan (1996) 1964/07-2019/12 0.22% 1.93%
Betting against beta BAB Frazzini and Pedersen (2014) 1964/07-2019/12 0.88% 3.34%
Cash flow to price CFP Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.30% 2.56%
Earnings to price ETP Basu (1983) 1964/07-2019/12 0.31% 2.63%
Liquidity LIP Pástor and Stambaugh (2003) 1968/01-2019/12 0.37% 3.34%
Long-term reversals LTR De Bondt and Thaler (1985) 1964/07-2019/12 0.20% 2.57%
Net share issues NSI Loughran and Ritter (1995) 1964/07-2019/12 0.27% 2.42%
Quality minus junk QMJ Asness, Frazzini, and Pedersen (2019) 1964/07-2019/12 0.42% 2.30%
Residual variance RVA Ang, Hodrick, Xing, and Zhang (2006) 1964/07-2019/12 0.24% 5.04%
Short-term reversals STR Jegadeesh (1990) 1964/07-2019/12 0.49% 3.15%
  • Note: This table provides the references, acronyms, sample periods, means, and standard deviations of the monthly returns for seven momentum factors and 15 U.S. factors examined in the empirical analysis.
4.2 Empirical Results
We first examine whether individual stock momentum and factor momentum contribute to the mean-variance efficient frontiers formed by the well-known Fama-French five factors. This assessment is accomplished by estimating the $\mathrm{MSS}$ for portfolios that combine one momentum factor with the five factors. The findings, as presented in Table 3, reveal consistent results across all factor combinations: the momentum factor, whether individual stock momentum or factor momentum, is consistently selected alongside the market, size, value, profitability, and investment factors. These results provide robust evidence that momentum factors significantly enhance the mean-variance efficient frontier established by the Fama-French five factors, corroborating the existing literature on stock return momentum; see Jegadeesh and Titman (1993); Carhart (1997); Rachev, Jašić, Stoyanov, and Fabozzi (2007); Ehsani and Linnainmaa (2022), among others.
Table 3: MSS for a Single Momentum Factor and the Fama-French Five Factors
Definition FTM PCM MOM IAM IDM ITM SRM
Momentum factor
Excess market return
Size
Value
Profitability
Investment
  • Note: This table presents the estimated MSS for portfolios comprising a single momentum factor and the Fama-French five factors. The momentum factor in the row corresponds to the acronyms listed in the respective columns. Detailed descriptions of the factors, their acronyms, and sampling periods are provided in Table 2. A check mark (✓) denotes that the corresponding asset is included in the estimated MSS. All analyses are performed at a significance level of 0.05.
Table 4: MSS for Stock Momentum, Factor Momentum, and the Fama-French Five Factors
Definition MON IAM IDM ITM SRM
Panel A
Stock momentum factor
Momentum in individual factors
Excess market return
Size
Value
Profitability
Investment
Panel B
Stock momentum factor
Momentum in PC factors 1–10
Excess market return
Size
Value
Profitability
Investment
  • Note: This table presents the estimated MSS for portfolios comprising a single stock momentum factor, a single factor momentum strategy, and the Fama-French five factors. Panel A evaluates sets that include individual momentum factors, while Panel B examines sets incorporating the first 10 principal component factors. In both Panels A and B, the stock momentum factor in the row corresponds to the acronyms listed in the respective columns. Detailed descriptions of the factors, their acronyms, and sampling periods are provided in Table 2. A check mark (✓) indicates that the corresponding asset is included in the estimated MSS. All analyses are conducted at a significance level of 0.05.
To distinguish the relative importance of individual stock momentum and factor momentum in shaping the efficient frontier, we next estimate the $\mathrm{MSS}$ for portfolios comprising one stock momentum strategy, one factor momentum strategy, and the Fama-French five factors. Table 4 reports the estimation results, highlighting three key findings. First, factor momentum is included in the estimated $\mathrm{MSS}$ for all ten cases, underscoring its critical role in constructing efficient frontiers. Second, although factor regressions suggest that individual stock momentum is largely explained by factor momentum (Ehsani and Linnainmaa, 2022), it nonetheless contributes to mean-variance efficient frontiers. Third, the market, size, profitability, and investment factors remain relevant in efficient frontiers when momentum in individual factors is present. However, the momentum in the first ten principal component factors (PCM) further subsumes the contribution of the size factor in efficient frontiers.
Our subsequent analysis evaluates the $\mathrm{MSS}$ for portfolios consisting solely of momentum factors, delineating which factors are most influential for the efficient frontier. Panel A of Table 5 details the estimated $\mathrm{MSS}$, consistently including the factor momentum strategy. However, only the PCM is selected when both factor momentum strategies are present. Additionally, various individual stock momentum factors such as standard stock momentum (MOM), industry-adjusted momentum (IAM), and Sharpe ratio momentum (SRM) significantly contribute to the efficient frontier when considering all related momentum factors. This finding contrasts with the findings of Ehsani and Linnainmaa (2022), which show that factor momentum fully accounts for stock return momentum. Instead, our results underscore the distinctive contribution of individual stock momentum in shaping mean-variance efficient frontiers.
Our estimation procedure provides more detailed insights into the relative importance of assets within the full set compared to the information presented in the table. To illustrate, Figure 2 displays plots of $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ along with their uniform confidence bands, for all momentum factors corresponding to the third column of Panel A in Table 5. The uniform confidence bands for $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ are centered at zero, with their upper bounds defined as $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,1}^{b}$ and $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,2}^{b}$, respectively. According to the estimation procedure, an asset is included in the $\mathrm{MSS}$ if, and only if, $|\hat{\alpha}_{i}|$ or $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both exceed their respective confidence bands.
From Figure 2, we observe that among the four momentum factors included in the estimated MSS, PCM is the most significant one. This is evident as both its $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ are substantially different from zero. For the other three momentum factors in the estimated MSS, their inclusion is driven by $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ being significantly different from zero, even though their $|\hat{\alpha}_{i}|$ values remain within the uniform confidence bands. Based on the statistic $M_{i,T}$, the ranking of momentum factors in the estimated MSS is as follows: PCM, IAM, SRM, and MOM.
Finally, we analyze the $\mathrm{MSS}$ for a comprehensive pooled portfolio comprising momentum factors, the Fama-French five factors, and ten additional well-established factors, as detailed in Panel B of Table 2. The estimation results are provided in Panel B of Table 5. To offer a graphical representation and further evaluate the relative importance of the factors included in the estimated MSS, we present plots of $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ along with their uniform confidence bands, for each factor in the full set in Figure 3, which corresponds to the results shown in the third column of Panel B in Table 5.
Table 5: MSS for Momentum Factors and Other Well-Established Factors
Definition FTM PCM FTM+PCM
Panel A
Momentum in individual factors N.I.
Momentum in PC factors 1–10 N.I.
Standard momentum
Ind.-adjusted momentum
Industry momentum
Intermediate momentum
Sharpe ratio momentum
Panel B
Momentum in individual factors N.I.
Momentum in PC factors 1–10 N.I.
Standard momentum
Ind.-adjusted momentum
Industry momentum
Intermediate momentum
Sharpe ratio momentum
Excess market return
Size
Value
Profitability
Investment
Accruals
Betting against beta
Cash flow to price
Earnings to price
Liquidity
Long-term reversals
Net share issues
Quality minus junk
Residual variance
Short-term reversals
  • Note: This table presents the estimated MSS for portfolios that include momentum factors and other factors. Panel A includes all momentum factors, while Panel B extends the analysis to portfolios comprising all momentum factors, the Fama-French five factors, and ten additional well-established factors. “N.I.” denotes that the corresponding factor is not included in the full set. Detailed descriptions of the factors, their acronyms, and sampling periods are provided in Table 2. A check mark (✓) indicates that the corresponding factor is included in the estimated MSS. All analyses are performed at a significance level of 0.05.
Figure 2: Graphical Illustration of the Two-step MSS Estimation Procedure
Note: This figure illustrates the estimation procedure for the $\mathrm{MSS}$, as analyzed in the last column of Panel A in Table 5. The assets are ordered from left to right based on their $M_{i,T}$ values, arranged in descending order (largest to smallest). The blue points above the $x$-axis represent the values of $|\hat{\alpha}_{i}|$, while the green points below the $x$-axis represent the values of $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$. The shaded regions above and below the $x$-axis indicate the 95% uniform confidence bands for $|\hat{\alpha}_{i}|$ and $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$, respectively. An asset is included in the $\mathrm{MSS}$ if, and only if, $|\hat{\alpha}_{i}|$, $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$, or both exceed their respective 95% uniform confidence bands. In such cases, the corresponding lines are highlighted in red.
From Panel B of Table 5, we observe that the size of the estimated $\mathrm{MSS}$ is considerably smaller than the total number of factors, highlighting the effectiveness of the proposed estimation procedure in reducing the number of assets required to replicate the mean-variance efficient frontier. Additionally, the estimated $\mathrm{MSS}$ consistently includes momentum factors across all three cases. Notably, PCM subsumes FTM when both are present in the full set. Moreover, IAM and MOM consistently contribute to the efficient frontier, whereas SRM, which appears in the estimated $\mathrm{MSS}$ in several cases in Panel A, is excluded in Panel B due to the presence of new factors in the full set. Furthermore, the estimated $\mathrm{MSS}$ also includes excess market return (MKT), size (SMB), accruals (ACC), betting against beta (BAB), quality minus junk (QMJ), and short-term reversal (STR), underscoring their significance in mean-variance analysis.
Figure 3: Graphical Illustration of the Two-step MSS Estimation Procedure
Note: This figure illustrates the estimation procedure for the MSS, as analyzed in the last column of Panel B in Table 5. The assets are ordered from left to right based on their $M_{i,T}$ values, arranged in descending order (largest to smallest). The blue points above the $x$-axis represent the values of $|\hat{\alpha}_{i}|$, while the green points below the $x$-axis depict the values of $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$. The shaded regions above and below the $x$-axis denote the 95% uniform confidence bands for $|\hat{\alpha}_{i}|$ and $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$, respectively. An asset is included in the MSS if, and only if, $|\hat{\alpha}_{i}|$, $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$, or both exceed their respective 95% uniform confidence bands. In such cases, the corresponding lines are highlighted in red.
From Figure 3, we observe that PCM, along with MKT and STR, are included in the estimated $\mathrm{MSS}$ because both their $|\hat{\alpha}_{i}|$ and $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values are significantly different from zero. These factors emerge as the most important assets in the full set in terms of their contributions to mean-variance efficiency. Two factors, BAB and QMJ, are selected due to their $|\hat{\alpha}_{i}|$ values being significantly above zero, even though their $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values fall within the uniform confidence bands. The remaining assets in the estimated $\mathrm{MSS}$—MOM, IAM, SMB, ACC, and RVA—are included because their $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values are significantly different from zero. Based on the statistic $M_{i,T}$, the ranking of factors in the estimated $\mathrm{MSS}$ is as follows: PCM, ACC, QMJ, SMB, MKT, IAM, MOM, STR, RVA, and BAB.
In summary, our analyses in this empirical application reaffirm the significant role of factor momentum in asset pricing. Factor momentum is consistently included in the $\mathrm{MSS}$ across all portfolios that incorporate individual stock momentum, factor momentum strategies, and other well-established factors. Notably, only the momentum in the first ten principal components contributes to the efficient frontier when coexisting with individual factor momentum. These findings align closely with those reported in Ehsani and Linnainmaa (2022). Additionally, certain individual stock momentum factors, such as the standard momentum factor and the industry-adjusted momentum factor, also enhance the construction of the mean-variance efficient frontier alongside factor momentum. Our method not only estimates the $\mathrm{MSS}$, but also provides valuable insights into the sources of each asset’s contribution to mean-variance efficiency and their relative importance within the full set of assets.
5 Conclusion
This paper proposes a method for estimating the $\mathrm{MSS}$, the smallest subset of assets capable of replicating the mean-variance efficient frontier of the full asset set. We derive the identification conditions for the $\mathrm{MSS}$ and propose an estimation and inference procedure based on these conditions. Under some regularity assumptions, we theoretically demonstrate that the proposed $\mathrm{MSS}$ estimator covers the true $\mathrm{MSS}$ wpa1, and converges to the true $\mathrm{MSS}$ with a probability reaching any pre-specified confidence level. A comprehensive set of Monte Carlo simulations shows that the procedure performs well in finite samples when the sample size is sufficiently large.
The proposed estimation and inference procedure is applied to analyze the $\mathrm{MSS}$ for a collection of individual stock momentum and factor momentum strategies, alongside other well-known factors. Our empirical study highlights the significant role of factor momentum from the perspective of mean-variance analysis, as it is consistently included in the estimated $\mathrm{MSS}$, when coexisting with individual stock momentum factors. This analysis also provides new insights into the relative importance of various individual momentum strategies and factors.
The asymptotic properties of our method are established under the assumption that the size of the full asset set, denoted as $d$ in the paper, is fixed. An interesting avenue for future research is to generalize the asymptotic framework to allow $d$ to grow with the sample size. Additionally, estimating the $\mathrm{MSS}$ in high-dimensional settings, where $d$ may exceed the sample size, presents a compelling research direction. The identification conditions established in this paper provide a guidance for designing valid estimation procedures in high-dimensional scenarios. We leave these extensions and related questions for future investigation.
Appendix
Appendix A.1 Proofs of the Main Results
This section presents the conditions required to establish the main results of the paper, namely Lemma 1, Theorems 1 and 2, along with their proofs. We begin by outlining the key conditions.
Assumption 1.
(i) $\left\{R_{t}\right\}_{t}$ is strictly stationary and strong mixing with mixing coefficient $\alpha_{j}$ satisfying $\alpha_{j}\leq Ca^{j}$ for some $a\in(0,1)$; (ii) $\mathrm{E}[\left|R_{i,t}\right|^{r}]\leq C$ for some constant $r>4$; (iii) the eigenvalues of $\mathrm{Var}(R_{t})$ are bounded away from zero; (iv) the dimension of $R_{t}$, denoted as $d$, is fixed.
Assumption 1 specifies conditions on the dimension, dependence structure, moment bounds, and variance-covariance matrix of the vector return process $\left\{R_{t}\right\}_{t}$. Specifically, Assumptions 1(i, ii, iv) are useful for establishing the consistency and the joint asymptotic normality of the estimators $(\hat{\theta}_{i})_{i\leq d}$. Meanwhile, Assumptions 1(ii, iii) ensure that the variance of $R_{t}$ is finite and non-singular, thereby guaranteeing that the mean-variance efficient frontier of $R_{t}$ is well-defined, not trivially reducible, and that the sufficient conditions of Lemma 1 are satisfied.
Assumption 2.
(i) the eigenvalues of $\Omega_{d,i}\equiv\lim_{T\rightarrow\infty}\mathrm{Var}(T^{-1/2}\sum_{t\leq T}e_{i,t})$are bounded away from zero; (ii) the bandwidth $\ell$ satisfies $\log(T)^{2}(\ell T^{-1/2+1/\tilde{r}}+\ell^{-1})=o(1)$ for some constant $\tilde{r}\in(4,r)$; (iii) $\gamma_{T}^{-1}T^{-1/8}=O(1)$.
Assumption 2 provides sufficient conditions to establish the validity of the bootstrap procedure for the estimation and inference of the MSS. Assumption 2(i) ensures that the asymptotic variance-covariance matrix of the estimator $\hat{\theta}_{i}$ is nonsingular for any $i\leq d$. Assumption 2(ii) imposes restrictions on the bandwidth $\ell$ used in the MBB procedure, specifically requiring that $\ell$ grows slower than $\log(T)^{-2}T^{1/2-1/\tilde{r}}$ but faster than $\log(T)^{2}$. Lastly, Assumption 2(iii) provides an upper bound on $\gamma_{T}$, which is satisfied by our recommendation, $\gamma_{T}=\min\{p/2,$ $0.05/\log(\log(T))\}$.
We now present the proofs of Lemma 1, Theorems 1 and 2. Throughout this section, $C$ denotes a positive constant that is independent of $i$, $t$ and $T$, and may vary from line to line.
Proof of Lemma  1. We begin by observing that $\Sigma^{-1}1_{d}(1_{d}^{\top}\Sigma^{-1}1_{d})^{-1}$ and $\Sigma^{-1}\mu(1_{d}^{\top}\Sigma^{-1}\mu)^{-1}$ represent the global minimum variance portfolio and the tangency portfolio (with zero risk-free rate), respectively. The claim of the lemma is proved in two separate cases.
First, consider the case that $\mu\neq a1_{d}$ for any $a\in\mathbb{R}$. In this case, the Markowitz algorithm (see, e.g., Theorem 3.1 in Constantinides and Malliaris (1995)) shows that the minimum variance portfolio $x_{e,\mu_{p}}$ with a target mean return $\mu_{p}$has a unique solution:
$x_{e,\mu_{p}}=\omega_{\mu_{p}}\frac{\Sigma^{-1}\mu}{1_{d}^{\top}\Sigma^{-1}\mu}+(1-\omega_{\mu_{p}})\frac{\Sigma^{-1}1_{d}}{1_{d}^{\top}\Sigma^{-1}1_{d}},$ (A.1)
where
$\omega_{\mu_{p}}\equiv\frac{(1_{d}^{\top}\Sigma^{-1}\mu)1_{d}^{\top}\Sigma^{-1}(\mu_{p}1_{d}-\mu\mathbf{)}}{(\mu^{\top}\Sigma^{-1}\mu)(1_{d}^{\top}\Sigma^{-1}1_{d}\mathbf{)-}(1_{d}^{\top}\Sigma^{-1}\mu\mathbf{)}^{2}}.$
The efficient portfolio frontier of $R$ is traced by $x_{e,\mu_{p}}$ as $\mu_{p}$ varies over $\mu_{p}\geq\mu^{\top}\Sigma^{-1}1_{d}(1_{d}^{\top}\Sigma^{-1}1_{d})^{-1}$. Therefore, the assets with non-zero weights in either $\Sigma^{-1}1_{d}(1_{d}^{\top}\Sigma^{-1}1_{d})^{-1}$ or $\Sigma^{-1}\mu(1_{d}^{\top}\Sigma^{-1}\mu)^{-1}$, or both, constitute the $\mathrm{MSS}$, as they form the minimal set of assets required to construct both these portfolios, which together span the efficient frontier of $R$. This implies:
$\mathrm{MSS}=\mathrm{Supp}_{\Sigma^{-1}\mu}\cup\mathrm{Supp}_{\Sigma^{-1}1_{d}}.$ (A.2)
Since $\Sigma^{-1}1_{d}$ and $\Sigma^{-1}\mu$ are both well-defined and non-zero, the existence and uniqueness of the $\mathrm{MSS}$ follow from (A.2). From Lemma A1, we have
$\mathrm{Supp}_{\Sigma^{-1}\mu}=\mathrm{Supp}_{(\alpha_{i})_{i\leq d}}\text{ \ \ and \ \ }\mathrm{Supp}_{\Sigma^{-1}1_{d}}=\mathrm{Supp}_{(1_{d-1}^{\top}\beta_{i}-1)_{i\leq d}},$ (A.3)
which, along with (A.2), establishes that the $\mathrm{MSS}$ satisfies (2.2).
In the case that $\mu=a1_{d}$ for some $a\in\mathbb{R}$, the efficient portfolio frontier of $R$ shrinks to $\Sigma^{-1}1_{d}(1_{d}^{\top}\Sigma^{-1}1_{d})^{-1}$. In this case
$\mathrm{MSS}=\mathrm{Supp}_{\Sigma^{-1}1_{d}}=\mathrm{Supp}_{\Sigma^{-1}\mu}\cup\mathrm{Supp}_{\Sigma^{-1}1_{d}},$ (A.4)
where the second equality holds because $\mathrm{Supp}_{\Sigma^{-1}\mu}$ is either empty (when $a=0$) or identical to $\mathrm{Supp}_{\Sigma^{-1}1_{d}}$ (when $a\neq 0$). Since $\Sigma^{-1}1_{d}$ is well-defined and non-zero, (A.4) confirms that the $\mathrm{MSS}$ exists and is unique. Combining the results from (A.3) and (A.4), we deduce that the $\mathrm{MSS}$ satisfies (2.2). $Q.E.D.$
Proof of Theorem  1. Let $\theta_{i,1}^{0}\equiv 0$ and $\theta_{i,2}^{0}\equiv 1$. By the triangle inequality
$\min_{i\in\mathrm{MSS}}M_{i,T}\geq\min_{i\in\mathrm{MSS}}\max_{j=1,2}\frac{\left|T^{1/2}(\theta_{i,j}-\theta_{i,j}^{0})\right|}{\hat{s}_{i,j}}-\max_{i\in\mathrm{MSS}}\max_{j=1,2}\frac{\left|T^{1/2}(\hat{\theta}_{i,j}-\theta_{i,j})\right|}{\hat{s}_{i,j}}.$ (A.5)
From Lemma A3, it follows that
$\max_{i\in\mathrm{MSS}}\max_{j=1,2}\frac{\left|T^{1/2}(\hat{\theta}_{i,j}-\theta_{i,j})\right|}{\hat{s}_{i,j}}=O_{p}(1).$ (A.6)
Let $Q\equiv\mathrm{E}[\tilde{R}_{t}\tilde{R}_{t}^{\top}]$ where $\tilde{R}_{t}\equiv(1,(R_{i,t})_{i\leq d}^{\top})^{\top}$. The matrix $Q$ can be expressed as:
$Q=\left(\begin{array}[]{cc}1&\mathbf{0}_{1\times d}\\ \mathrm{E}[R_{t}]&I_{d}\end{array}\right)\left(\begin{array}[]{cc}1&\mathbf{0}_{1\times d}\\ \mathbf{0}_{d\times 1}&\mathrm{Var}(R_{t})\end{array}\right)\left(\begin{array}[]{cc}1&\mathrm{E}[R_{t}^{\top}]\\ \mathbf{0}_{d\times 1}&I_{d}\end{array}\right).$
Therefore, by Assumptions  1(ii, iii), we have
$C^{-1}\leq\lambda_{\min}(Q)\leq\lambda_{\max}(Q)\leq C,$ (A.7)
where $\lambda_{\min}(Q)$ and $\lambda_{\max}(Q)$ denote the smallest and largest eigenvalues of $Q$, respectively. From (A.7), it further follows that
$\min_{i\leq d}\sigma_{\varepsilon_{i}}^{2}\geq\lambda_{\min}(Q)>C.$ (A.8)
Combininig the results in (A.7) and (A.8) with Lemma A2, we have
$\min_{i\in\mathrm{MSS}}\max_{j=1,2}\frac{\left|T^{1/2}(\theta_{i,j}-\theta_{i,j}^{0})\right|}{\hat{s}_{i,j}}$$\geq$ $(C^{-1}-o_{p}(1))\min_{i\in\mathrm{MSS}}\max_{j=1,2}\frac{\left|T^{1/2}(\theta_{i,j}-\theta_{i,j}^{0})\right|}{\left\|A_{j,\cdot}\right\|}$ (A.9)
$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2},$
which, along with (2.16), (A.5) and (A.6), implies that
$\min_{i\in\mathrm{MSS}}M_{i,T}\geq(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}.$ (A.10)
Since $\overline{cv}_{1-p,T}^{b}=O_{p}(1)$ by Lemma A4(i), from (A.10) we obtain:
$\min_{i\in\mathrm{MSS}}M_{i,T}-\overline{cv}_{1-p,T}^{b}$$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}-O_{p}(1)$ (A.11)
$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2},$
where the second inequality is by (2.16). From (A.11), it follows that:
$\mathrm{P}\left(\mathrm{MSS}\subset\overline{\mathrm{MSS}}_{p}\right)=\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(C^{-1}-o_{p}(1)>0\right)\geq 1-o(1),$ (A.12)
which shows the first claim of the theorem. [^11] Since the $\mathrm{MSS}$ is nonempty, (A.12) implies that $\overline{\mathrm{MSS}}_{p}$ is nonempty wpa1. Therefore, the finite sample adjustment in Step 7 of the algorithm in Subsection 2.2, when applied to construct $\overline{\mathrm{MSS}}_{p}$, is asymptotically negligible.
For any $p\in(0,1)$, let $\overline{cv}_{1-p}^{u}$ denote the ($1-p$)th quantile of $\max_{i\leq d}\tilde{M}_{i}$. Then by Lemma A5, there exists a positive sequence $\delta_{T}=o(1)$ such that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\delta_{T}}^{u}\right)-o(1).$ (A.13)
Therefore for any $\varepsilon\in(0,1-p)$, we have
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)-o(1),$ (A.14)
for sufficiently large $T$. By Lemma A3, it follows that
$\lim_{T\rightarrow\infty}\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)=\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)\geq\mathrm{P}\left(\max_{i\leq d}\tilde{M}_{i}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right),$ (A.15)
where the inequality holds because $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq\max_{i\leq d}\tilde{M}_{i}$. This, combined with (A.14) and the definition of $\overline{cv}_{1-p-\varepsilon}^{u}$ shows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq 1-p-\varepsilon-o(1).$ (A.16)
Collecting the results from (A.12) and (A.16), we obtain:
$\mathrm{P}\left(\mathrm{MSS}=\overline{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)$ (A.17)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)+\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)-1$
$\geq$ $1-p-\varepsilon-o(1).$
Since $\varepsilon$ is arbitrary, the claim of the lemma follows from (A.17). $Q.E.D.$
Proof of Theorem  2. Since $cv_{1-p,T}^{b}\leq\overline{cv}_{1-p,T}^{b}$ by construction, from Lemma A4(i) and the similar arguments for showing (A.11), we obtain
$\min_{i\in\mathrm{MSS}}M_{i,T}-cv_{1-p,T}^{b}$$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}-O_{p}(1)$
$\geq$ $(C^{-1}-o_{p}(1))\gamma_{T}^{-1/2},$
where the last inequality is by (2.19). Hence
$\mathrm{P}\left(\mathrm{MSS}\subset\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.18)
$\geq$ $\mathrm{P}\left((C^{-1}-o_{p}(1))\gamma_{T}^{-1/2}>0\right)\geq 1-o(1),$
which shows the first claim of the lemma. [^12] This also shows that the finite sample adjustment in Step 7 of the algorithm in Subsection 2.2 is asymptotically negligible.
To show the second claim, we first note that by Lemma A6,
$cv_{1-p,T}^{b}=cv_{1-p,T}^{b}(\overline{\mathrm{MSS}}_{\gamma_{T}})=cv_{1-p,T}^{b}(\mathrm{MSS})$ (A.19)
wpa1. Therefore, by Lemma A5,
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)$$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS})\right)+o(1)$ (A.20)
$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p+\delta_{T}}^{u}(\mathrm{MSS})\right)+o(1),$
where $\delta_{T}=o(1)$ is a positive sequence. From (A.20), it follows that for any $\varepsilon\in(0,p)$, we have
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p+\varepsilon}^{u}(\mathrm{MSS})\right)+o(1),$
for sufficiently large $T$, which together with Lemma A3 shows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon+o(1).$ (A.21)
Therefore,
$\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.22)
$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon+o(1).$
To derive the lower bound for $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})$, we apply (A.19) to obtain
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS})\right)-o(1).$ (A.23)
This, along with Lemma A5, shows that for any $\varepsilon^{\prime}\in(0,1-p)$
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)$$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\delta_{T}}^{u}(\mathrm{MSS})\right)-o(1)$ (A.24)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right)-o(1),$
for sufficiently large $T$. By Lemma A3, we have
$\lim_{T\rightarrow\infty}\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right)=\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right).$ (A.25)
From (A.24), (A.25) and the definition of $cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})$, it follows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\geq 1-p-\varepsilon^{\prime}-o(1),$
which together with (A.18) shows that
$\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.26)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)+\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)-1$
$\geq$ $1-p-\varepsilon^{\prime}-o(1).$
Since $\varepsilon$ and $\varepsilon^{\prime}$ are arbitrary, the second claim of the lemma follows from (A.22) and (A.26). $Q.E.D.$
Appendix A.2 Auxillary Lemmas
This section presents some auxiliary results that are utilized in proving Lemma 1, Theorems 1 and 2 in the previous section. For any $i=1,\ldots,d$, let $\mu_{i}\equiv$ $\mathrm{E}[R_{i}]$, $\mu_{-i}\equiv$ $\mathrm{E}[R_{-i}]$, $\sigma_{i}^{2}\equiv\mathrm{Var}(R_{i})$ and $\Sigma_{-i}\equiv\mathrm{Var}(R_{-i})$. Define $\tilde{\sigma}_{i}^{2}\equiv\sigma_{i}^{2}-\Gamma_{i,-i}\Sigma_{-i}^{-1}\Gamma_{i,-i}^{\top}$, where $\Gamma_{i,-i}\equiv\mathrm{Cov}(R_{i},R_{-i})$.
Lemma A1.
Suppose that $\Sigma$ is finite and non-singular. For any $i=1,\ldots,d$, the regression coefficients in (2.1) satisfy $\alpha_{i}=\tilde{\sigma}_{i}^{2}\ell_{d,i}^{\top}\Sigma^{-1}\mu$ and $1_{d-1}^{\top}\beta_{i}-1=\tilde{\sigma}_{i}^{2}\ell_{d,i}^{\top}\Sigma^{-1}1_{d}$, where $\tilde{\sigma}_{i}^{2}>0$.
Lemma A2.
Under Assumption 1, we have $\max_{i\leq d}\max_{j=1,2}|\hat{s}_{i,j}^{2}/s_{i,j}^{2}-1|=O_{p}(T^{-1/2})$.
Lemma A3.
Under Assumption 1, we have for any nonempty subset $\mathcal{S}$ of $\mathcal{M}$:
$\max_{i\in\mathcal{S}}\max_{j=1,2}\frac{\left|T^{1/2}(\hat{\theta}_{i,j}-\theta_{i,j})\right|}{\hat{s}_{i,j}}\rightarrow_{d}\max_{i\in\mathcal{S}}\tilde{M}_{i},$
where $\tilde{M}_{i}\equiv\left\|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d}^{1/2}\mathcal{N}_{d}\right\|_{\infty}$ and $\Omega_{d}=\lim_{T\rightarrow\infty}\Omega_{d,T}$.
For the $t$-th observation $R_{t}^{b}$ in the bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}$, let $R_{-i,t}^{b}$ denotes its subvector excluding $R_{i,t}^{b}$. Define $\tilde{R}_{-i,t}^{b}\equiv(1,(R_{-i,t}^{b})^{\top})^{\top}$ and $\varepsilon_{i,t}^{b}\equiv R_{i,t}^{b}-\theta_{i}^{\top}\tilde{R}_{-i,t}^{b}$. For any proper subset $\mathcal{S}$of $\mathcal{M}$, including the empty set, let $cv_{1-p}^{u}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}\tilde{M}_{i}$, and $cv_{1-p,T}^{b}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}M_{i,T}^{b}$. Moreover, we define:
$\tilde{M}_{i,T}^{b}\equiv\max_{j=1,2}\frac{\left|T_{B}^{-1/2}\sum_{t\leq T_{B}}e_{i,j,t}^{b}\right|}{s_{i,j}},$
with $e_{i,j,t}^{b}\equiv A_{j,\cdot}Q_{-i}^{-1}(\tilde{R}_{-i,t}^{b}\varepsilon_{i,t}^{b}-\mathrm{E}^{\ast}[\tilde{R}_{-i,t}^{b}\varepsilon_{i,t}^{b}])$ and $\mathrm{E}^{\ast}[\cdot]$ denoting the expectation taken under the bootstrap distribution given the data.
Lemma A4.
Let $\overline{cv}_{1-p,T}^{b}\equiv cv_{1-p,T}^{b}(\varnothing)$. Under Assumptions 1 and 2, we have (i) $\overline{cv}_{1-p,T}^{b}=O_{p}(1)$ for any $p\in(0,1)$; and (ii) $\overline{cv}_{1-\gamma_{T},T}^{b}=O_{p}(\gamma_{T}^{-1/2})$.
Lemma A5.
Under Assumptions 1 and 2, there exists a positive sequence $\delta_{T}=o(1)$ such that for any $p\in(0,1)$,
[…]
Lemma A6.
Under Assumptions 1 and 2, we have $\mathrm{MSS}=\overline{\mathrm{MSS}}_{\gamma_{T}}$ wpa1.
Appendix A.3 Additional Simulation Results
Figure 4: Empirical Distribution of the Intersection Size Between the Estimated and True MSS
Note: This figure illustrates the empirical distribution of the size of the intersection between the estimated and true MSS for four different settings. The upper panel corresponds to the settings of $(K,N,T)=(1,7,120)$ and $(K,N,T)=(1,7,240)$, respectively, while the lower panel corresponds to the settings of $(K,N,T)=(7,1,120)$ and $(K,N,T)=(7,1,240)$. The significance level is set to 0.05.